OpenStudy (danysanchy97):

How many more people preferred WWCN than WANR? The graph represents the results of a survey in which 400 people were asked to identify the TV channel on which they preferred to watch theevening news. A. 164 B. 236 C. 68 D. 76

rebeccaxhawaii (rebeccaxhawaii):

whats the context

OpenStudy (danysanchy97):

WWCN-41% WANR-22%

OpenStudy (danysanchy97):

@Dani_Rose @Daniee_Bruhh

OpenStudy (dani_rose):

i have no idea sorry

OpenStudy (danysanchy97):

its ok

OpenStudy (daniee_bruhh):

I might know how to do this lol give me a few minutes

OpenStudy (danysanchy97):

@imqwerty

OpenStudy (danysanchy97):

ok thanks @Daniee_Bruhh

OpenStudy (danysanchy97):

sorry i frogot to add that part to the question

OpenStudy (dani_rose):

100% = 400 people. WWCN = 41% of 400 people 41/100 X 400 = 164 people WANR = 22% 400 people 22/100 X 400 = 88 people WCLM = 24% of 400 people 24/100 X 400 = 96 people WKOD = 13% of 400 people 13/100 X 400 = 52 people ```````````````````````` People who prefered WCLM=96 Subtract: 164 - 88 = 76

OpenStudy (danysanchy97):



OpenStudy (dani_rose):

lol i had this question

OpenStudy (danysanchy97):

lol thanks

OpenStudy (dani_rose):

lol you're welcome <3

OpenStudy (daniee_bruhh):

lol No need for me to explain now XD Good Luck on your test anyways love! :3

OpenStudy (danysanchy97):

thanks @Daniee_Bruhh

OpenStudy (daniee_bruhh):

No Problem love! :D

OpenStudy (danysanchy97):

wait @Dani_Rose why did yo put WKOD and WCLM in there?

OpenStudy (dani_rose):

Last time i had the problem i need ed to include them for other questions lol I thought you would need them

OpenStudy (danysanchy97):

oh ok thanks